Okay here is my question...
I know how to find the hybridisation of (say),PCl5 molecule.It has 40 valence electrons and so divided by 8 it gives sp3d hybridisation and it has a trigonal bipyramidal geometry but my question is hoq do i find the hybrisation of a particular atom like Phosphorous in Pcl5 ????

Help me yaar-aieee givin me nightmares !!!
God bless repliers please!!!

Grade:12

2 Answers

Umesh Kumar
13 years ago

If the atom consists of single bond then it is sp3 hyberdised atom, if double then it is sp2, if triple then sp

vikas askiitian expert
13 years ago

dear geoffery ,   hybridisation of PCl5 is hybridisation of central atom (P) of this molecule ...

hybridisation of N in NH3 is SP3 & O in H2O is SP3 ...

 

short trick for hybridisation =>

n =[ number of electron in valence shell + number of monovalent atoms attached(like H,Cl etc) - charge on molecule]/2

for n = 2 - hybridisation is SP ,

n =3 hybridisation is SP2 ,

for n = 4 - SP3  & so on  .........

 

eg=> NH3 ..... valence electrons = 5 , monovalent atoms = 3 , charge = 0 so

                    n = [5+3-0]/2 = 4 ,  hybridisation is SP3 

 

eg=> NH4+, n = 5+4-1/2 = 4     ,   SP3

 

eq=> CO2 , n = 4/2 = 2 = SP                           (here in this molecule monovalent atoms are 0)